### v1

# Hands-on STEM Learning Experiences using Digital Technologies

Abstract
The facilitation of STEM education can be enhanced by the provision of opportunities for learners to gain a better understanding of science through the utilization of tangible and visual examples. The objective of this work is to present an account of our experiences and activities carried out in Italian schools with this novel approach. The selection of projects and experiences discussed –in which students develop a range of core competencies such as collaboration, creativity, critical thinking, experimentation, prototyping, communication and problem-solving; include tangible complex 3D printed structures, large micro-controller board replicas and the visualization of wind dynamics and tiny invisible elementary particles among others. These hands-on experiences demonstrate the benefits on the use of digital fabrication technologies implemented within a FabLab for STEM learning.
STEM Education; School Support; Inclusion; Fabrication Laboratory
I Introduction
The present approach to develop hands-on science, technology, engineering, and mathematics (STEM) learning experiences using digital technologies, is motivated by the 2030 UN agenda for sustainable development –which provides a global framework for progress and sustainable development. In order to achieve these goals, it is essential to implement innovative strategies that will lead to improvements in STEM education and the creation of opportunities to more scholars. In particular, the Sustainable Development Goal 4 of the UN 2030 Agenda aims to “Ensure inclusive and equitable quality education and promote lifelong learning opportunities for all” [1].
In the last decade on-line education has transformed society to new levels in the way people teach and learn [2]. Additionally, a number of other significant digital innovations are currently in place which collectively aim to facilitate the achievement of the 2030 UN goals. Examples are the use of 3D printing for the production of Mathematical and Physics educational structures, the application of micro-controllers and tiny computers for sensor-based experiments [3]. These are becoming increasingly popular, which are helping to stimulate scholars’ curiosity and encourage a more profound understanding of specific subjects. Thus, learning is today facilitated under an authentic, engaging, personal context in which scholars engage in reflection, imagination, comprehension and comparison of solutions through the use of specific, tangible examples. The implementation of digitally-controlled technologies for tangible and visual STEM learning is enabling scholars to access the highest quality education.
In this work we describe some of our experiences applying a hands-on and visual approach to STEM education, developed in our Scientific Fabrication Laboratory (FabLab) which is equipped with the necessary facilities to build specific devices for that purpose. The specifications of these devices are openly shared on-line. We argue that placing learning literally in the hands of the younger generations is a more effective approach, specially when it does not pose any special technical requirements on the teachers’ side.
Furthermore, our objective is also to develop techniques and processes that will enable new generations of students to participate fully in the educational process, in a manner that is inclusive of individuals with physical disabilities [4].
II Previous Works
Some twenty-first century skills are difficult to teach only by traditional methods, i.e., without the use of highly sophisticated information and communications technologies (ICT). However, further studies on the design, creation and implementation of new instruments in which the teaching-learning process can be more personalized is still needed –see a review in [5]. The use of ICT has been recurrent in the last decade, specially due to the massive diffusion of portable PCs, tablets and smartphones [2]. Nevertheless, alternative digital technologies available at FabLabs can be employed to assist on the development of innovative pedagogical approaches that are accessible to all students and that facilitate their understanding on the relevance of scientific concepts [3].
FabLabs enable the transformation of an idea into a tangible object, making it a potential pedagogical tool in STEM education. The establishment of FabLabs in Academia can play a relevant role as a resource for the creation of pedagogical objects and devices to support sustainable development within the educational context. Fablabs facilitate the implementation of tangible and visual STEM learning, thereby opening up new possibilities. The following sections will present a selection of our projects and experiences, in which students develop a range of core competencies, including collaboration, creativity, critical thinking, experimentation, prototyping, communication, and problem-solving. A discussion to better understand the influence of Makerspaces and FabLabs on today’s innovation processes can be found in [6]. STEM education opportunities for everyone via the implementation of Makerspaces and FabLabs in vulnerable and marginal communities is discussed in [7]. On the other hand, the potential aspects of FabLabs in relation to traditional books as a pedagogical tool, and how teachers perceive these tools for teaching Chemistry to their students, have been investigated in [8].
The evidence gathered from these implementations lend support to STEM learning practices via FabLabs as a rewarding and meaningful experience for all their participants, and also help to enlarge community adoption of educational digital tools in surrounding areas by the word of mouth.
III Some STEM Prototypes
The following section presents a selection of tangible and visual examples of STEM devices that can be reproduced at any FabLab. FabLabs are typically equipped with low-cost, state-of-the-art and versatile computer-controlled rapid prototyping tools including 3D printers and scanners, laser engraving and CNC cutting machines driven by open source software. FabLabs provide a working space for invention, creativity and resourcefulness and they are strongly rooted in a local community of makers. STEM students can learn more effectively through practice and be more productive through experience with the prototypes produced in these FabLabs.
The aim of our FabLab is also to disseminate scientific knowledge and to provide opportunities for people with disabilities to engage with science. The results presented here are the consequence of experiences derived from research activities conducted at the ICTP Scientific Fabrication Laboratory of Trieste, Italy, over the past few years. These experiences were conducted as hands-on, one-off, learning experiences specially in Italian primary and secondary schools, and in activities during events such as Science Picnics and Maker Faires.
III.1 3D STEM Objects
The current state of low-cost 3D printing technologies is such that they have reached a level of maturity that allows for their unlimited potential to be realized in the field of STEM education. The capacity to reproduce 3D objects of varying complexity, ranging from intricate mathematical structures up to sophisticated biological specimens, has not only benefited schools and higher education institutions in countries with limited scientific infrastructure to sustain their STEM education programs, but it has also facilitated the rapid grow of digital STEM content around the world.
Typically, 3D printing with FDM (Fused Deposition Modeling) technology is conducted using spools of thermo-plastic filament like polylactic acid (PLA) – an environmentally friendly material derived from corn starch. A number of free or proprietary tools for 3D modeling and printing are available on the net, facilitating the work required to design and print the objects.
Figure 1: Left: Examples of complex 3D printed mathematical structures. Right: Scale model of a giant extinct elephantid genus Mammuthus.
III.2 Large Micro-controller Board Replica
Microcontrollers, such as the popular Arduino Uno, can be employed in a multitude of educational projects including robotics, the monitoring of thermodynamic variables via sensors (e.g., humidity, temperature, pressure, etc), RGB Led-based controllers, and GPS trackers among others. This technology is becoming increasingly accessible, offering a high degree of affordability and versatility. The introduction of microcontrollers to primary school students can be facilitated through the use of block-based coding.
A comprehensive examination of the use of microcontrollers in primary and secondary schools has revealed that the primary obstacle to their implementation is the limited scale of electronic components and the associated difficulty in handling them. This has prompted us to develop a larger version of the most commonly used microcontroller, Arduino Uno, together with a prototype board and some electronic components with magnetic connections. The larger (9:1) scale facilitates the handling of the attached components while maintaining the overall set’s ease of transportation. The scale model was constructed using a laser cutter and a combination of 3D printed components, connectors and pins.
The fully functioning giant replica in Fig.2 not only allows an instructor to explain and demonstrate how to use the microcontroller to students in a classroom setting, but it also benefits students with visual impairments or motor skill difficulties, enabling them to participate actively in the lectures.
Figure 2: Large micro-controller Arduino board replica.
III.3 DIY Lift Generator for STEM: Tubora
In order to teach wind dynamics and test which object does fly the best and what does not, a Do-it-Yourself (DIY) wind tube can be built using just simple elements such as a fan and a transparent plastic tube as shown in Fig.3. This system is designed to regulate the pull of air throughout the tube and, in addition, the tube can be configured to point at different angles.
Figure 3: Tubora: an invaluable resource for students engaged in STEM studies of aerodynamics or the mechanics of parachutes.
Placing the fan on the underside of the large tube, and inserting objects of varying materials, students can observe the different aerodynamic characteristics of each object. As they collect and analyze data to ascertain the mechanisms underlying the movement of objects in a forward or downward direction, or their retention in the air at varying angles of flow, they gain insight into the most intriguing aspects of natural physics, namely the phenomena associated with objects in an airflow, aerodynamics and variable drag forces.
The versatility of this exhibit is noteworthy. It allows children in the first years of primary school (aged 6-8) to construct paper objects in a creative manner, thereby determining the probability of them flying out of the tube. This provides students with the opportunity to experiment with different shapes, testing them each time, and to color them in an imaginative way. In addition to cones and paper airplanes, students can investigate the efficacy of twirling paper helicopters. This involves varying factors such as the type of paper, the addition of paperclips, the size of wings, and so forth.
Figure 4: STEM Engineering Activity: the runaway zoo. Examples of activity sheets.
they are presented with an object (specifically a small lightweight stuffed animal) that is too heavy to be lifted by the fan alone. Each group is provided with the same materials (paper, tulle, cotton and umbrella fabric, plastic bags, string, masking tape) and must design the most efficient solution to allow the “skydiver” to rise and escape from the tube. The students will have the opportunity to experiment with different approaches to the design of the parachute canopy. They will select the material they wish to use for the main canopy and then proceed through the engineering design process, which involves defining the problem, brainstorming potential solutions, creating prototypes and testing them in a wind tunnel. Each group will also be required to complete a record sheet, which will help them document their choices, the reasons behind them, and some background information on the plush animal and its need to escape. They are also required to draft sketches of the front and top views of their prototype to complete their task thus fulfilling the final step of the Engineering design process (document and share) –see Fig.4. This parachute challenge offers the opportunity to investigate a multitude of aspects related to flight and the effects of force on an object. It also provides an opportunity to examine the engineering design process, which is a systematic approach to problem-solving.
III.4 Highly Visual Cloud Chamber –Physics in Action!
A facility for studying tiny invisible elementary particles that enter our atmosphere from the outer space (cosmic rays) or that are generated by the decay of radioactive elements within the Earth itself, and then pass right through our bodies, is the well-known apparatus known as “Wilson’s cloud chamber”. This apparatus was invented in the early 1900s by Charles Wilson as one of the first particle detectors. The majority of these particles are generated by the decay of natural radon found in the air (itself a product of radioactive element’s decay in the ground) while a certain percentage, which varies with altitude, does originate from the cosmic radiation we receive from the outer space (primary and secondary cosmic rays).
Figure 5: Typical particles tracks in a cloud chamber device (Left). Full video available at https://www.youtube.com/watch?v=IwyLzqrRJbw Right: Public interest on the DIY cloud chamber.
To build a cloud chamber for educational purposes one needs a transparent box with a black bottom metal tray maintained at a temperature of approximately minus 30 degrees Celsius, accompanied by the presence of a stable cloud of minute droplets of isopropyl alcohol, typically generated by condensation from a liquid state. After a period of time, the top lid of the chamber, which has been soaked in warm alcohol, produces a cloud that reaches the cold metal bottom. When a charged particle passes through at a high speed it generates a visible track of larger droplets by a mechanism known as coalescence. The droplets of alcohol can then be detected due to their ability to diffuse the light emitted by the strips of LEDs positioned on the sides, as illustrated in Fig.5.
The particles that can be observed in a cloud chamber can be broadly classified into three categories: alpha particles, beta particles and muons. Alpha particles are composed of two protons and two neutrons. They have a short, fat track and are produced by the decay of radon. Beta particles are electrons (or anti-electrons, a.k.a. positrons), usually following long straight paths or zigzag paths, depending on the energy and momentum of the particles. Muons are particles emitted from the decay of secondary cosmic rays and can be observed as short vertical tracks. Further details on the construction and operation of a cloud chamber can be found in [9].
III.5 Galileo and Leonardo Mechanical Clocks
The first application of a pendulum in a clock used an escapement system is attributed to Galileo Galilei from his original drawings. Following the discovery of the isochronism of a pendulum by Galileo’s observation of a swinging lamp in Pisa Cathedral (dating back to 1600), a working gravity-powered clock can be made on a laser cutter with the addition of standard bearings and a few parts made from plywood. As shown in Fig.6 (Left), the design of a Galileo mechanical clock can be written using OpenSCAD –a free, fully customizable software.
Figure 6: OpenSCAD design of a Galileo mechanical clock (Left). Video demo available at: https://www.youtube.com/watch?v=a6B1bxgxAc4 Right: Leonardo Mechanical Clock. Demo video available at: https://www.youtube.com/shorts/g4hkIHu89kE
Galileo’s clock mechanism uses gravitational potential energy as its power source, and carries two arcs; a small lower arc that receives the impulse, while an upper wide arc lifts a detent. This model has been designed in collaboration with the “Amici dell’orologeria Pesarina” association, engaged since many years in the dissemination of knowledge concerning watchmaking, a tradition present in the area of Pesariis (a ancient village in the mountains of Friuli Venezia Giulia region, Italy) since 1600, and its dissemination in local schools. The request was to produce a model escapement that could be made cheaply and then to introduce a kit into secondary schools so that students could experience the mechanisms first hand. The files for this build can be dowloaded freely (https://www.thingiverse.com/thing:5549040) and can be further customized in OpenSCAD easily.
The Leonardo Da Vinci Clock depicted in Fig.6 (Right) is another wooden pendulum clock that can be constructed in a FabLab for educational purposes (https://www.thingiverse.com/thing:4328105). This clock model is powered by a horizontal rotary pendulum, and utilizes the weight contained in a canister and some gears to measure time. The canister has stored potential energy gained by raising it up against gravity. The potential energy is converted to kinetic energy allowing the gears to move their parts. By adjusting the weight and balance of the pendulum, the clock’s speed can then be controlled.
III.6 Giant Galton Board
A large Galton board allows students to visualize the preferred trajectories of falling (ping-pong) balls while being fed into the device from the top at arbitrary rates. Due to the differing initial conditions established by the participants, the gravitational force and the presence of ordered, fixed obstacles along their paths, the individual balls experience random elastic collisions as they descend. The falling balls positions will not concentrate at specific locations: Instead, they will be characterized by a distribution (with one or more peaks) resulting from the influence of the stochastic effects due to the presence, or absence, of the obstacles. The spatial distribution of the falling balls with homogeneous obstacles is observed to exhibit a symmetric bell-shaped (Gaussian) curve –see Fig.7.
Figure 7: Giant Galton board of descent ping-pong balls.
The observation of individual ball trajectories represents the dynamics of a single ensemble member, whereas the bell statistical distribution of the final balls positions represents the total ensemble average for the system. This device can serve as a learning platform for explaining weather forecast models, as they use the equations governing the atmosphere in conjunction with a stochastic algorithm called “Montecarlo” to generate a weather forecast (corresponding to the resulting distribution of balls). The model, as well as the device, exhibits a high degree of sensitivity to its initial conditions (of the dropped balls), where even minor initial perturbations give rise to divergent weather predictions. The presence or absence of the obstacles correspond to the various parameters of the underlying model, and they can be moved to “tune” the results.
III.7 Augmented Reality Sandbox
The Augmented Reality (AR) Sandbox is an educational hands-on exhibit developed by the University of California, Davis, USA. It combines a box filled with real sand, with virtual topography and simulated water. The apparatus is constructed using a Microsoft Kinect 3D camera, in addition to a video projector and a Linux PC to achieve extensive real-time computer simulations and visualization. The AR Sandbox allows the reconstruction of geographical models through the shaping of the sand. These models are then augmented by an elevation color map, topographic contour lines, and simulated liquid water.
Figure 8: The Sandbox unit (Left) which displays a mountain with a crater lake, surrounded by several smaller lakes (Right). This tool is an invaluable resource for educators and learners alike, facilitating the acquisition of knowledge and understanding in the domains of topography, geography, natural sciences and water flow.
The device, illustrated in Fig.8, has been employed for science communication projects involving experiments on geological hazards and volcano alerts. The system is useful for illustrating concepts related to modeling volcanoes, including the processes involved in volcano deformation and the study of magma reservoirs.
A variety of hands-on experiments can be conducted, such as placing a small, inflated balloon beneath the sand. Deflating the balloon can then demonstrate the effect of a collapsing volcano magmatic chamber. It is also possible to illustrate the impact of the rising sea level on coastal areas: for instance Easter Island –located in the southeastern Pacific Ocean, is highly vulnerable to the effects of climate change, and this can be visually demonstrated with great clarity by varying digitally the height of the sea level.
III.8 2D-to-3D Image Conversion via Artificial Intelligence Model
3D visualization has the potential to enhance the understanding of STEM among young students, thereby facilitating more effective learning. Virtual learning opens up new worlds in STEM education, particularly in the context of abstract concepts and the development of specific experiences [10]. Virtual learning can facilitate the exploration of the interior of a human cell, traverse planetary surfaces, or navigate subterranean magma chambers. These are examples of the ”intangible” world that can be experienced in virtual learning environments. In order for this to be achieved, it is necessary to recreate a realistic 3D environment. It would be even more beneficial if the viewer were able to perceive this content more naturally, i.e., without the need for special 3D glasses.
Figure 9: Outputs of ”altiro3D” algorithm to represent reality. Left: original single RGB image. Center: multiview synthesis of a single image carried out through artificial intelligence. Right: 3D display on a headset-free Lenticular 3D screen. Video demo available at: https://www.youtube.com/shorts/hJDVb2TzBr0
A visual 3D scene representation of a learning environment serves as a medium through which the integration of STEM learning may be effectively transformed into a tangible and meaningful experience. It has a tremendous potential to revolutionize education, yet it remains to be further explored. The free library ”altiro3D, which enables the representation of reality from a single 2D RGB image or flat video, has also been developed in our Scientific FabLab [11]. This software allows the synthesis of multiple N-viewpoints images or video to create a realistic (real-time) 3D immersive experience when displayed on a lenticular, free-view LCD screen as shown in Fig.9. altiro3D does not require heavy computing runtime and can support a wide range of application scenarios in education and science. The synthesis of multiple views from a single image is achieved through the integration of machine learning together with deep learning algorithms, and simple inpainting techniques, which map all image pixels.
IV Discussion
The current configuration of public primary and secondary schools in Italy presents certain challenges. The minimum number of pupils per class in primary schools is 15, with a maximum of 26 pupils permitted, except in the case of disabled pupils. In secondary schools, the minimum number of students per class is 18, with a maximum of 27 students permitted [12].
A further challenge is posed by the increasing number of students from other countries: in the 2021/2022 academic year, 10.6% of students in Italy were from other countries, with the figure in Friuli Venezia Giulia (FVG) Region reaching 15.4% and 22.7% in the province of Gorizia [13]. A considerable number of educational institutions are currently undergoing renewal works, with classes being temporarily relocated to smaller classrooms or to alternative buildings that have been converted into schools. The challenges of working in such circumstances are considerable, particularly when the classes are overcrowded and a significant proportion of students lack proficiency in the local language. STEM curricula are highly engaging for students, particularly those with language difficulties. The hands-on nature of many STEM subjects enables all students to participate fully, regardless of their linguistic abilities. This phenomenon is further evidenced when students with disabilities are present in the classroom. According to our experiences, STEM activities can be modified to accommodate the specific needs of these students, with the objective of overcoming physical limitations and addressing mental impairments. During a lesson in which a class with a blind student was being taught, the possibility of combining a buzzer with the LED outcome of the experiment proved invaluable in fully involving the student in the lesson.
On another occasion, a student in a wheelchair was able to participate completely in the activity by having all the educational material relocated to the correct height for him to reach. In another instance, a student with an autism spectrum disorder had the time needed to fully appreciate the effects of air thrust on a small parachute, while his fellow students were engaged in the collection of data and the preparation of a subsequent demonstration.
While curricula teachers interact with students throughout the academic year, structuring their lessons over a long time span, STEM laboratories are often conducted by external educators with the aid of specific materials and teaching aids, and must accommodate the needs of the school timetable. The principal objective of these laboratories is to stimulate interest among students and provide the curricula teachers with insights into alternative pedagogical approaches to technical subjects. The lessons provided can then be further developed during the school year.
The experiences of one of the authors (G.F.) on providing one-off, hands-on learning activities in primary and secondary schools over the past decade has been predominantly focused on the delivery of engaging, 1.5-3 hour long sessions that engage the entire class, while also providing all the necessary materials. This approach has the dual benefit of relieving teachers of the burden of providing materials while also allowing for more focused and effective learning.
Figure 10: Students STEM activities for Italian Schools from the FVG Region carried out at the Scientific FabLab of ICTP.
Students currently engaged in primary and secondary education have experienced the challenges of the Covid-19 lockdown and a prolonged period of distance learning, which has resulted in significant difficulties in working in groups and reduced dexterity. This kind of learning hands-on experience using digital technologies, can be extremely useful in addressing both these problems. In our experience, dividing the class into four groups is optimal, as it allows for more focused and effective learning while allowing a single educator (such as the situation usually presents) to monitor the progress of each group, ensuring that all groups remain on track while providing additional challenges to the more advanced group.
While working with smaller groups or a one-to-one approach would also be interesting, the limitations posed by the need to engage the whole class and have a single educator limit the possibilities, resulting in an ideal configuration with four groups of students, forcing them to work as teams, exchanging ideas and comparing their different approaches, allowing for a more focused and effective learning. Furthermore, this configuration has enabled teachers to examine the interactions between their students in a group setting, offering valuable insights. The decision regarding the composition of groups is always a challenging one, with students frequently attempting to maintain proximity to their acquaintances, thereby creating groups that are not optimally balanced. When working with primary school pupils the involvement of teachers, who possess a comprehensive understanding of the students and are capable of dividing them (maintaining the separation of more challenging students and creating more balanced groups) can be invaluable in this regard. In the context of working with secondary school students, it is possible to allow them to choose their own group, thereby giving them the opportunity to assume this responsibility.
In most cases, the curricula teachers participate, assisting students with more difficulties and providing important context with references to lessons already given on similar topics. Frequently, they choose to discuss in advance the topics covered during such activities, stimulating students’ curiosity before the laboratory, creating an optimal learning environment. The class is often requested to provide feedback on the STEM topics presented, which may take the form of a simple drawing or collaborative poster for primary school students, or a summary or more creative medium such as a video, comic, rap song, newscast or other form for secondary students. The variety of media employed and the creativity displayed by classes that have previously encountered disciplinary problems during the laboratory is often noteworthy.
V Conclusions
STEM students can greatly benefit when receiving training and education based on tangible and visual learning. The prototypes presented in this work as examples allows to put concrete experiences in their hands. The aforementioned examples have been generated utilizing digital technologies that are readily accessible in any FabLab worldwide. We have discussed and assessed some of our experiences and activities done in Italian Schools using such devices. For example, by touching complex 3D structures, visualizing cosmic rays, and building the needed blocks to learn micro-controllers in class.
We have verified how hands-on STEM learning emphasizes students potential and opens a new path to present them with a smooth connection between the understanding of abstract scientific phenomena and a deeper, more comprehensive learning experience. This novel approach in action is appealing to scholars and can sustain new generations of scholars in an inclusive way facilitating the quality of STEM education for all.
Acknowledgement
We thank Sara Sossi, Erika Ronchin and Marco Baruzzo who have helped with our ICTP SciFabLab STEM activities in the last few years. Sincere gratitude is also due to Fondazione Pietro Pittini (FPP) for partnering with us in many of these projects.